Consider an electron which has been  accelerated from rest through a potential difference of 500kV. Find a) Its kinetic energy, b.) its rest energy, c.) its total energy, d.) its mass and e.) its speed.


1
Expert's answer
2022-02-13T12:15:28-0500

a)

"E_k=eU=0.5~MeV,"

b)

"E_0=mc^2=0 .512~MeV,"

c)

"E=E_k+E_0=1~MeV,"

d)

"m=\frac E{c^2}=1.8\cdot10^{-30}~kg,"

e)

"v=c\sqrt{1-\frac{m_0^2}{m^2}}=0.87c."